Question6 10 pts Suppose there are two people, with their willingness to pay for a good given by WTP1 10- 2Q and WTP2 5- Q/2. If the good in question is a private good, what is the aggregate demand for these two people? 。WTP-10-2Q for 0 Q < 2.5 and WTP 6-20/5 otherwise. O WTP = 10-2Q for 0 s Q < 5 and WTP = 6-20/5 otherwise. 。WTP-15-SQ/2 for 0 Q < 2.5 and WTP-5-Q/2 otherwise. ·WTP = 15-5Q/2 for 0 Q < 5 and WTP = 5-Q/2 otherwise. Question 7 10 pts Suppose there are two people, with their willingness to pay for a good given by WTP1 10 - 2Q and WTP2-5- Q/2. If the good in question is a public good, what is the aggregate demand for these two people? 。WTP-10-2Q for 0 Q < 2.5 and WTP = 6-20/5 otherwise. O WTP 10- 2Q for Os Q5 and WTP 6- 2Q/5 otherwise. WTP = 15-50/2 for 0 Q < 2.5 and WTP = 5-Q/2 otherwise. 。WTP-15-SQ/2 for 0 Q < 5 and WTP-5-Q/2 otherwise.
Answer #1

6) For private good we add the original demand function / WTP function

2Q = 10 - WTP1 and Q/2 = 5 - WTP2

Q = 5 - 0.5WTP and Q = 10 - 2WTP

Add them to get private demand at Q = 15 - 2.5WTP or WTP = 6 - 2Q/5 for Q greater than 5 (Second option)

7) For public good just add the two inverse demand functions

WTP = 10 - 2Q + 5 - 0.5Q or WTP = 15 - 2.5Q. Since minimum of Q is 5, last option is correct.